OpenStudy (anonymous):

what is the slope of a line parallel to 2x-3y=7

OpenStudy (anonymous):

get it in the form y = mx +b

OpenStudy (anonymous):

y=2/3x-7/3

OpenStudy (anonymous):

yep so the slope of your parallel line will have the sam slope as this line

OpenStudy (anonymous):

okay so does that me the formula is just the same because that is what i have but i am just checking to see if i am right

OpenStudy (anonymous):

the formula for a parallel line to your line y=2/3x-7/3 can be in the form y=2/3x + a where a is any number positive or negative

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

but its best to chose a as something other than what your original one was a = -7/3 just so the lines arent overlapping

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

one more question if the the formula isy=3/5x+45 wouldnt 3/5 be the x-intercept?

OpenStudy (anonymous):

you can find the x intercept by setting y =0 then solve for x

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

your welcome